Two tiny, spherical water drops, with identical charges of -8.73 × 10-16 C, have a center-to-center separation of 1.35 cm. (a) What is the magnitude of the electrostatic force acting between them? (b) How many excess electrons are on each drop, giving it its charge imbalance?
Answer #1

Force acting between tiny water drops is

Then,

F=K q1q2/r^2

Put values..

F=(9*10^9)(8.73*10^-16)^2 /(0.0135m)^2

F=3.7636*10^-17 N Answer

Part B)

No. Of electrons is

=-8.73*10^-16 C / 1.6*10^-19 C

=5456 electrons Answer

I hope help you .